# Noisy-Pair Robust Representation Alignment for Positive-Unlabeled Learning

### v1

@@ Abstract
Positive-Unlabeled (PU) learning aims to train a binary classifier (positive vs. negative) where only limited positive data and abundant unlabeled data are available. While widely applicable, state-of-the-art PU learning methods substantially underperform their supervised counterparts on complex datasets, especially without auxiliary negatives or pre-estimated parameters (e.g., a 14.26% gap on CIFAR-100 dataset). We identify the primary bottleneck as the challenge of learning discriminative representations under unreliable supervision. To tackle this challenge, we propose NcPU, a non-contrastive PU learning framework that requires no auxiliary information. NcPU combines a noisy-pair robust supervised non-contrastive loss (NoiSNCL), which aligns intra-class representations despite unreliable supervision, with a phantom label disambiguation (PLD) scheme that supplies conservative negative supervision via regret-based label updates. Theoretically, NoiSNCL and PLD can iteratively benefit each other from the perspective of the Expectation-Maximization framework. Empirically, extensive experiments demonstrate that: (1) NoiSNCL enables simple PU methods to achieve competitive performance; and (2) NcPU achieves substantial improvements over state-of-the-art PU methods across diverse datasets, including challenging datasets on post-disaster building damage mapping, highlighting its promise for real-world applications. Code: Code will be open-sourced after review.
@@ 6 Related Works
Contrastive and Non-contrastive Representation Learning. Self-supervised learning has recently made significant progress in acquiring high-quality representations. Depending on the use of negative pairs, self-supervised learning methods are broadly categorized into contrastive and non-contrastive learning: contrastive losses encourage embeddings of similar images to be aligned (alignment) while pushing apart those of different images (uniformity) (He et al., 2020; Wang & Isola, 2020; Huang et al., 2023a; Lu et al., 2023); non-contrastive losses only align embeddings of similar images Grill et al. (2020); Chen & He (2021); Guo et al. (2025). Beyond self-supervised settings, SupCon was proposed to extend contrastive learning by incorporating full supervision (Khosla et al., 2020). Recent works actively explore the application of contrastive learning to weakly supervised tasks (Li et al., 2022; Huang et al., 2023b; Zhang et al., 2024; Wang et al., 2022; 2024; Yuan et al., 2025). However, most of these methods rely on supervised contrastive learning with pseudo labels, which suffer from the problem of noisy pairs. The WSC framework (Zhou et al., 2025), a graph-theoretic weakly supervised contrastive learning method, introduces continuous semantic similarity; however, it requires pre-estimated parameters as input. In contrast, NcPU, which builds on non-contrastive learning, is robust to noisy pairs and does not rely on such requirements.
@@ 7 Conclusion
This paper identifies the challenge of learning discriminative representations as the key bottleneck in PU learning. Theoretical analysis demonstrates that noisy pairs dominate representation learning optimization, while the proposed NoiSNCL and PLD form an EM-inspired framework that iteratively benefits each other. Extensive evaluations on both benchmark datasets and remote sensing building damage mapping tasks validate the effectiveness of the proposed NcPU. Applications of binary classification with only positive data annotated can benefit from this study, and we expect future research to extend this framework beyond image classification. Beyond PU learning, the framework holds promise for broader applications in more weakly supervised learning scenarios.
@@
An anonymous code repository (Code will be open-sourced after review) is provided to facilitate the reproduction of the experimental results. Theoretical results, including assumptions and complete proofs, are presented in detail in the main text and the appendix. All datasets used in this work are publicly available and have been widely adopted in the research community.
@@ Appendix G More Experimental Results
The results of precision (P), recall (R), and area under receiver operating characteristic curve (AUC) are presented in Table 5-Table 6. NcPU achieves a better trade-off between P and R.
@@ Appendix H More Ablation Analyses
$\tilde{\mathcal{L}}_{\text{r}}$ significatly enhances the performance of PU learning methods. Apart from Table 3, we also conducted another set of experiments to test the robustness of $\tilde{\mathcal{L}}_{\text{r}}$ with respect to fewer training data and higher $\pi_{p}$. CIFAR-100 is used as the experimental dataset, with the division of positive and negative classes consistent with the main experiments, and ResNet-18 serves as the backbone. To highlight the effectiveness of $\tilde{\mathcal{L}}_{\text{r}}$, the $\pi_{p}$ is set to 0.6, with 1000 positive training samples and 20000 unlabeled training samples. Optimization is performed using SGD with a momentum coefficient of 0.9 and a cosine annealing learning rate scheduler. The initial learning rate is set to 0.001. The $w_{\text{r}}$ is set to 100. The pair-construction strategy is generalized by introducing a similarity threshold ($\tau$), which enables more flexible construction of pairs:
@@
Appendix I The Use of Large Language Models

### v2

@@ Abstract
Positive-Unlabeled (PU) learning aims to train a binary classifier (positive vs. negative) where only limited positive data and abundant unlabeled data are available. While widely applicable, state-of-the-art PU learning methods substantially underperform their supervised counterparts on complex datasets, especially without auxiliary negatives or pre-estimated parameters (e.g., a 14.26% gap on CIFAR-100 dataset). We identify the primary bottleneck as the challenge of learning discriminative representations under unreliable supervision. To tackle this challenge, we propose NcPU, a non-contrastive PU learning framework that requires no auxiliary information. NcPU combines a noisy-pair robust supervised non-contrastive loss (NoiSNCL), which aligns intra-class representations despite unreliable supervision, with a phantom label disambiguation (PLD) scheme that supplies conservative negative supervision via regret-based label updates. Theoretically, NoiSNCL and PLD can iteratively benefit each other from the perspective of the Expectation-Maximization framework. Empirically, extensive experiments demonstrate that: (1) NoiSNCL enables simple PU methods to achieve competitive performance; and (2) NcPU achieves substantial improvements over state-of-the-art PU methods across diverse datasets, including challenging datasets on post-disaster building damage mapping, highlighting its promise for real-world applications. Code: https://github.com/Hengwei-Zhao96/NcPU.
11footnotetext: *Corresponding author.
@@ 5.3 Ablation Studies and Analyses
Figure 6: OA curve of NcPU during the training process. NcPU exhibits stable training performance over a prolonged training process.
Analysis of training stability. While the gradients of $\tilde{\mathcal{L}}_{\text{r}}$ do not vanish and include the term $\frac{1}{\sqrt{1-\bm{\tilde{q}}_{i}^{\top}\bm{\tilde{k}}_{j}}}$, they do not lead to significant overfitting or training instability: (1) Benefiting from the asymmetric architecture and random data augmentation, $\bm{\tilde{q}}_{i}\neq\bm{\tilde{k}}_{j}$ is ensured, which can enable the numerical stability of the gradient; (2) As $\bm{\tilde{q}}_{i}^{\top}\bm{\tilde{k}}_{j}$ approaches 1, the gradient magnitude becomes $\frac{2}{\left\lVert\bm{q}_{i}\right\rVert_{2}^{2}}$ instead of infinity, within a finite number of training iterations, $\tilde{\mathcal{L}}_{\text{r}}$ does not lead to notable overfitting or training instability; (3) The training stability can also be empirically verified (Figure 6): Taking the results on CIFAR-10 dataset as an example, NcPU has achieved promising performance at the 400th epoch. After extended training for a longer period, the results of NcPU does not exhibit overfitting or instability; (4) The numerical stability can be improved by constraining the value of $\bm{\tilde{q}}_{i}^{\top}\bm{\tilde{k}}_{j}$ (e.g., $10^{-4},1-10^{-4}]$).
@@ 6 Related Works
Contrastive and Non-contrastive Representation Learning. Self-supervised learning has recently made significant progress in acquiring high-quality representations. Depending on the use of negative pairs, self-supervised learning methods are broadly categorized into contrastive and non-contrastive learning: contrastive losses encourage embeddings of similar images to be aligned (alignment) while pushing apart those of different images (uniformity) (He et al., 2020; Wang and Isola, 2020; Huang et al., 2023a; Lu et al., 2023); non-contrastive losses only align embeddings of similar images (Grill et al., 2020; Chen and He, 2021; Guo et al., 2025). Beyond self-supervised settings, SupCon was proposed to extend contrastive learning by incorporating full supervision (Khosla et al., 2020). Recent works actively explore the application of contrastive learning to weakly supervised tasks (Li et al., 2022; Huang et al., 2023b; Zhang et al., 2024; Wang et al., 2022; 2024; Yuan et al., 2025). However, most of these methods rely on supervised contrastive learning with pseudo labels, which suffer from the problem of noisy pairs. The CoTAP Loss (Wen et al., 2025) is proposed to align the representations of semantically similar objects in self-supervised dense representation learning, and the negative effect of noises is alleviated by assigning higher weights to sample pairs with top scores. The WSC framework (Zhou et al., 2025), a graph-theoretic weakly supervised contrastive learning method, introduces continuous semantic similarity; however, it requires pre-estimated parameters as input. In contrast, NcPU, which builds on non-contrastive learning, is robust to noisy pairs and does not rely on such requirements.
@@ 7 Conclusion
This paper identifies the challenge of learning discriminative representations as the key bottleneck in PU learning. Theoretical analysis demonstrates that noisy pairs dominate representation learning optimization, while the proposed NoiSNCL and PLD form an EM-inspired framework that iteratively benefits each other. Extensive evaluations on both benchmark datasets and remote sensing building damage mapping tasks validate the effectiveness of the proposed NcPU. Applications of binary classification with only positive data annotated can benefit from this study, and we expect future research to extend this framework beyond image classification. Masked image modeling will be incorporated to harness more powerful backbone frameworks. Beyond PU learning, the framework holds promise for broader applications in more weakly supervised learning scenarios.
@@
A code repository (https://github.com/Hengwei-Zhao96/NcPU) is provided to facilitate the reproduction of the experimental results. Theoretical results, including assumptions and complete proofs, are presented in detail in the main text and the appendix. All datasets used in this work are publicly available and have been widely adopted in the research community.
@@ Appendix G More Experimental Results
The results of precision (P), recall (R), and area under receiver operating characteristic curve (AUC) are presented in Table 5-Table LABEL:tab_results_stl10/abcd. NcPU achieves a better trade-off between P and R.
@@ Appendix H More Ablation Analyses
$\tilde{\mathcal{L}}_{\text{r}}$ significatly enhances the performance of PU learning methods. Apart from Table 3, we also conducted another set of experiments to analyze the contribution of representation quality on PU learning and test the robustness of $\tilde{\mathcal{L}}_{\text{r}}$ with respect to fewer training data and higher $\pi_{p}$. CIFAR-100 is used as the experimental dataset, with the division of positive and negative classes consistent with the main experiments, and ResNet-18 serves as the backbone. To highlight the effectiveness of $\tilde{\mathcal{L}}_{\text{r}}$, the $\pi_{p}$ is set to 0.6, with 1000 positive training samples and 20000 unlabeled training samples. Optimization is performed using SGD with a momentum coefficient of 0.9 and a cosine annealing learning rate scheduler. The initial learning rate is set to 0.001. The $w_{\text{r}}$ is set to 100. The pair-construction strategy is generalized by introducing a similarity threshold ($\tau$), which enables more flexible construction of pairs:
@@
Analysis on Imbalanced Datasets. NcPU has been validated under more challenging conditions: class imbalance (Table 11) and distribution imbalance (Su et al., 2021; Zhao et al., 2023b) (Table 11). WSC was selected as the comparison method due to its best performance without any auxiliary information. In class imbalance scenarios, the number of positive samples is significantly smaller than that of unlabeled samples. The imbalance ratio (IR) is used to quantify the degree of class imbalance, defined as the ratio of the number of unlabeled samples to positive samples. IR can be increased by reducing the number of positive samples (all other settings remain consistent with the main experiments). Results in Table 11 demonstrate that NcPU is robust to class imbalance. Recent studies (Su et al., 2021; Zhao et al., 2023b) have indicated that distribution imbalance is another critical challenge for PU learning tasks: a low $\pi_{p}$ exerts adverse effects on classifier training. In the experiments of Table 11, 20000 unlabeled samples were used, and distribution imbalance was simulated by decreasing the $\pi_{p}$ (all other settings are consistent with the main experiments). Results in Table 11 show that NcPU is robust to distribution imbalance.
Table 10: Class imbalance results on CIFAR-10 dataset.
Method IR: 80 IR: 53.33
OA F1 OA F1
WSC 81.46±3.0 71.31±6.4 89.70±0.7 86.59±1.4
NcPU(ours) 93.70±4.0 91.34±6.1 96.44±0.9 95.42±1.2
Table 11: Distribution imbalance results on CIFAR-10 dataset.
Method $\pi_{p}$: 0.05 $\pi_{p}$: 0.1
OA F1 OA F1
WSC 90.97±0.3 88.48±0.4 90.72±0.3 88.36±0.4
NcPU(ours) 93.57±0.5 91.36±0.7 93.88±0.4 91.81±0.6
@@
Appendix I Analysis on Computational Overhead
The comparison of computational overhead between NcPU and other existing PU learning methods in both training and inference phases is presented in Table 12. Among these methods, nnPU is a classic PU learning algorithm without a representation learning module, while LaGAM, WSC, and NcPU all incorporate representation learning modules: (1) In the training phase, compared with nnPU, the PU learning algorithms with representation learning modules exhibit increased training time per epoch, but there is no significant difference in the per-epoch training time between NcPU and the other algorithms with representation learning modules; (2) In the inference phase, NcPU and other methods only utilize a single classification network for inference, thus NcPU achieves the same inference speed and computational complexity to the other methods; (3) Compared with other methods with representation learning modules, NcPU does not show a significant increase in per-epoch training time, yet it delivers a substantial improvement in OA and F1.
Table 12: Computational overhead of different methods on CIFAR-10 dataset.
Method Training Phase Inference Phase Task Performance
Training Time/Epoch (s) Batch Inference
(ms/256 samples) GFLOPs OA F1
nnPU 5.06 12.92 0.56 87.29±0.5 83.71±0.6
LaGAM 14.79 95.78±0.5 94.90±0.6
WSC 16.00 90.55±0.3 87.92±0.8
NcPU 14.84 97.36±0.1 96.67±0.2
Appendix J The Use of Large Language Models